A company is trying to decide which of two types of tires to buy for their trucks. They would like to adopt to brand C unless there is some evidence that brand D is better. An experiment was conducted where 16 tires from each brand were used. The tires were run under similar conditions until they wore out. The results are:

 

Brand C:  1 = 40,000 kms., s1 = 5,400 kms.

Brand D:  2  =38,000 kms., s2 =3,200 kms.

What conclusions can be drawn?


1
Expert's answer
2021-09-27T13:23:18-0400

"n_1=n_2=16 \\

\bar{x_1}=40000 \\

s_1 = 5400 \\

\bar{x_2}= 38000 \\

s_2=3200 \\

H_0: \mu_1 < \mu_2 \\

H_1: \mu_ 1> \mu_2"

When n1 and n2 are small and "\sigma_1" and "\sigma_2" are unknown, we have to use the two-sample t test for independent random samples from two normal populations having the same unknown variance.

Test-statistic:

"t= \frac{\bar{x_1} - \bar{x_2}}{s_p \sqrt{(1/n_1) + (1/n_2)}} \\

s^2_p= \frac{(n_1-1)s^2_1 +(n_2-1)s^2_2}{n_1+n_2-2} \\

s^2_p= \frac{(16-1)(5400)^2 +(16-1)(3200)^2}{16+16-2} \\

= \frac{4.374 \times 10^8 + 1.536 \times 10^8}{30} \\

= 0.197 \times 10^8 \\

s_p=0.4438 \times 10^4 \\

= 4438.4 \\

t= \frac{40000 -38000}{4438.4 \sqrt{(1/16) + (1/16)}} \\

= \frac{2000}{1569.2} \\

= 1.274"

Reject H0 if "t\u2265t_{crit}"

Let use α=0.05

For one-tail test and degree of freedom "df = n_1+n_2-2= 30 \; t_{crit} = 1.697"

"t=1.274 < t_{crit}= 1.697"

Accept H0 at 0.05 significance level.

Therefore, there is enough evidence to claim that the population mean μ1 is less than the population mean μ2. Brand D is better than C.